Recipe has 2 cups of flour with 1.5 cups liquid. If I want to increase liquid to 2 cups, how much flour would I use?
Answer by math_tutor2020(3817)


(2 cups flour)/(1.5 cups liquid) = (x cups flour)/(2 cups liquid)
2/1.5 = x/2
2*2 = 1.5*x
4 = 1.5x
1.5x = 4
x = 4/1.5
x = 4/(3/2)
x = 4*(2/3)
x = 8/3
x = 2 & 2/3

You need to use 2 & 2/3 cup of flour.